OpenStudy (anonymous):

Express the complex number in trigonometric form. -6i

OpenStudy (sidsiddhartha):

first let \[-6i=a+ib\] so what's ur a and b?

OpenStudy (anonymous):

I don't know. I don't understand this.

OpenStudy (sidsiddhartha):

look here \[a=real part \] and \[b=imaginary part\] so a=0 and b=-6 do u get it?

OpenStudy (anonymous):

Yes. so -6 = 0 + -6i?

OpenStudy (sidsiddhartha):

no we are just comparing -6i with a+ib and we get b=-6 as a=0 all right yes or no?

OpenStudy (anonymous):

okay. yes

OpenStudy (sidsiddhartha):

now if u wanna Express any complex number in trigonometric form then u have write it like this \[z= a + ib = -6i = R(\cosΘ + isinΘ)\] where \[R = √(a²+b²) = √(0+-6²) = 6\] and \[Θ = \arctan(b/a) = \arctan(b/0) = \arctan(\infty)=π/2 \] so far all good?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Does that mean it is 6( cos 90 +i sin 90)?

OpenStudy (sidsiddhartha):

yup very good

OpenStudy (anonymous):

YAY! Thank you!

OpenStudy (sidsiddhartha):

now a very essential thing is left look coordinates of (a,b) is=(0,-6) is'nt it

OpenStudy (sidsiddhartha):

so u have to and pi=180 degree more with each angle so it will become\[6[\cos(180+ 90) +i \sin (180+90)]=6(\cos270+isin270)\] that's ur answer